What is the coordination number of K4(FeCN6)? I know the answer but I don't know how we will get it. Please explain me in detail?

Grade:

2 Answers

Arun
4 years ago

EAN means metal atomic no. Minus oxidation no.of metal atoms plus 2× coordination number,which is equal to next atomic number.

  • In k4(fe(cn)6) , central atom Fe. I..e, 26.
  • Oxidation no.=4(x(-1)6) =2
  • So Fe is +2
  • Coordination no=2×6 =12
  • Then EAN =26–2×12 =36
  • 36 is stable configuration.
Vikas TU
4 years ago
Dear student 
can be calculated as follows,
4(+1) + x + 6(-1) = 0
Hence x = +2.
K3[Fe(CN)6] , we know that Potassium, K, has an oxidation number of +1
 
 
 The oxidation number of CN= -1, so 6*CN has an oxidation number of -6 so therefore
 
 Fe - 6 = -3
 
 Fe = -3 + 6 = +3
 
 The oxidation number of Fe is +3